Problem 2. (10pts.) Find the general solution to the system of differential equations y = 2 - 2 z = 5y + 42
Answer #1

Y=2-7 z = 54+42 Dy +z =2 - -59 +(-4)2 = 0 ☺ x5 ©xD → 50y + 5z = 10 -5 Dy + D(D-4)2=0 (2 - 4D+5)z=10 Auxillary eqn j of a ma_> y=5688t Get - 4 Gest_46et 23 =centagets 2= cest toeta